OpenStudy (lollylau):

Maxima and Minima: Find the smallest distance of from a point from x^2 to 10,0.

OpenStudy (anonymous):

lagrange?

OpenStudy (lollylau):

Round off the the nearest tenths.

OpenStudy (lollylau):

I'm still thinking -.-

OpenStudy (paxpolaris):

minimum of:\[\sqrt{(x-10)^2+(x^2-0)^2}\]

OpenStudy (lollylau):

I got an answer that looks really wrong.

OpenStudy (lollylau):

Anyone got an answer?

OpenStudy (turingtest):

patience...

OpenStudy (lollylau):

>.<

OpenStudy (anonymous):

grad F=<2(x-10),2y> grad G=<-2x,1> 2(x-10)= -2x lambda 2x-20=-2x lambda 2y= lambda y-x^2=0 http://www.wolframalpha.com/input/?i=Solve%5B2x-20%3D-2x%20*a%2C%202y%3Da%2C%20y-x%5E2%3D0%5D&t=crmtb01 x=1.61 y=2.60

OpenStudy (across):

PaxPolaris is correct; the distance formula between two points \((x_1,y_1)\) and \((x_2,y_2)\) is \(\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\). Since the function \(f(x)=x^2\) can be expressed as an ordered pair by \((x,x^2)\), and you're given the point \((10,0)\), you can substitute those values instead.\[\]

OpenStudy (lollylau):

Thats what I got...

OpenStudy (anonymous):

what did you get?

OpenStudy (lollylau):

your ans

OpenStudy (anonymous):

I used lagrange multiplier ; which is useful for finding min/max while staying on curve/surface

OpenStudy (lollylau):

?

OpenStudy (lollylau):

I consider this problem solved.

OpenStudy (paxpolaris):

you can just forget the square root get the minimum of the square of the distance